The Ferrell Transportation Company uses a responsibility reporting system to measure the performance of its three investment centers: Planes, Taxis, and Limos. Segment performance is measured using a system of responsibility reports and return on investment calculations. The allocation of resources within the company and the segment managers’ bonuses are based in part on the results shown in these reports.

Recently, the company was the victim of a computer virus that deleted portions of the company’s accounting records. This was discovered when the current period’s responsibility reports were being prepared. The printout of the actual operating results appeared as follows.

Determine the missing pieces of information below. (Round intermediate calculations and final answers to 0 decimal places, e.g. 1,255.)

Planes

Taxis

Limos

Service revenue

$

$498,500

$

Variable costs

5,502,400

297,900

Contribution margin

250,900

522,340

Controllable fixed costs

1,499,600

Controllable margin

83,000

277,740

Average operating assets

24,960,500

1,543,000

Return on investment

13

%

7

% %
Answer #1

Determine the missing pieces of information below. (Round intermediate calculations and final answers to 0 decimal places, e.g. 1,255.)

Planes

Taxis

Limos

Service revenue

$10246865

$498,500

$820240

Variable costs

5,502,400

498500-250900 = 247600

297,900

Contribution margin 4744465

250,900

522,340

Controllable fixed costs

1,499,600

250900-83000 = 167900 244600
Controllable margin 24960500*13% = 3244865

83,000

277,740

Average operating assets

24,960,500

83000*100/7 = 1185714

1,543,000

Return on investment

13

%

7

% 18 %